OpenStudy (anonymous):

For which equation does the point (20,60) lie on the graph ? Explain i. C=15+2.50n ii. C=50+1.75n iii. C=30+1.50n

OpenStudy (anonymous):

i dont understand how to do it

hartnn (hartnn):

@minotti Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) the point (20,60) lie on the graph means it satisfies the equation of the graph. So, you have to plug in the point (n,C) = (20,60) in all the choices and check which one equation gets satisfied by this point.

OpenStudy (anonymous):

can you show me how to figure it out.

hartnn (hartnn):

ok, i will plug in that point in the first equation/choice and see whether the point satisfies the equation of the graph. i. C=15+2.5n (n,C)=(20,60) put n=20 C = 15 + 2.5*20 = 15 +50 = 65 but C needed to be 60 , so C=15+2.5n is not the equation of required graph. You check for ii. and iii.

OpenStudy (anonymous):

ok hold on trying to fiqure it out

OpenStudy (anonymous):

c=50+1.75n (n,c)=(20,60) n=20 c=50+1.75 20=50+75=125

hartnn (hartnn):

1.75*20 =... ?

OpenStudy (anonymous):

do i mupliy it

hartnn (hartnn):

yes.

OpenStudy (anonymous):

ok 1.75*20=3.500

hartnn (hartnn):

1.75*20 = 35 c=50+1.75n (n,c)=(20,60) n=20 c=50+1.75 20=50+35=85 but C needed to be 60 , so C=50+1.75n is not the equation of required graph. now check iii.

OpenStudy (anonymous):

c=30+1.50n (n,c)=(20,60) n=20 c=30=1.50 20=30+30=60

hartnn (hartnn):

yes ! so we got C=60 as required. so, the graph is C=30+1.50n

OpenStudy (anonymous):

yes i understand it thank you for helping me

hartnn (hartnn):

welcome ^_^